### v1

# Uncertainty-Weighted Image-Event Multimodal Fusion for Video Anomaly Detection

Abstract
Most existing video anomaly detectors rely solely on RGB frames, which lack the temporal resolution needed to capture abrupt or transient motion cues—key indicators of anomalous events. To address this limitation, we propose Image-Event Fusion for Video Anomaly Detection (IEF-VAD), a framework that synthesizes event representations directly from RGB videos and fuses them with image features through a principled, uncertainty-aware process. The system (i) models heavy-tailed sensor noise with a Student’s-t likelihood, deriving value-level inverse-variance weights via a Laplace approximation; (ii) applies Kalman-style frame-wise updates to balance modalities over time; and (iii) iteratively refines the fused latent state to erase residual cross-modal noise. Without any dedicated event sensor or frame-level labels, IEF-VAD sets a new state of the art across multiple real-world anomaly detection benchmarks. These findings highlight the utility of synthetic event representations in emphasizing motion cues that are often underrepresented in RGB frames, enabling accurate and robust video understanding across diverse applications without requiring dedicated event sensors. Code and models are available at https://github.com/EavnJeong/IEF-VAD.
1 Introduction
Recent advances in deep learning have led to significant progress in multimodal data analysis [72, 38, 23, 54, 37, 26, 1], enabling systems to effectively integrate diverse sensory inputs for complex tasks such as video anomaly detection [17, 76, 60, 32, 14]. Traditionally, research in this area has predominantly focused on leveraging static image information, which provides rich spatial context. However, dynamic event data—capturing subtle motion dynamics, abrupt changes, and transient temporal patterns—remains largely underexplored. Event data inherently offers high temporal resolution and can capture fleeting anomalies that static images might miss [19, 55, 63, 5]. Although the event modality has recently been integrated into multimodal learning frameworks [31, 82, 71], its practical potential remains underrealized, primarily due to the scarcity of real-world event datasets. In this work, we introduce to use synthetic event data extracted from videos to overcome this limitation.
The challenge of effectively combining heterogeneous data sources—static image information and dynamic event cues—under conditions of uncertainty remains a critical open problem [22, 77, 20]. Recent transformer-based fusion methods for video anomaly detection [16, 66, 83, 85, 12, 22] still focus on the rich spatial details in images and, as a result, under-utilize the transient yet crucial temporal cues in event streams. Because multimodal attention mechanisms naturally gravitate toward the more expressive modality, this bias suppresses the complementary information that other inputs could provide [53, 62, 41]. These limitations highlight the need for a fusion strategy that can balance modality contributions under uncertainty.
To address these challenges, we propose an approach that explicitly fuses the event modality with image data through an uncertainty-aware framework. Our Uncertainty-Weighted Image-Event Fusion framework for Video Anomaly Detection (IEF-VAD) balances the two modalities by assigning inverse-variance weights derived from Bayesian uncertainty estimates. By modelling each modality’s latent features alongside their predictive variance, IEF-VAD down-weights less reliable signals and prevents the image stream—richer in spatial content but often dominant—from overshadowing the temporally informative event cues. Concretely, we capture the heavy-tailed sensor noise of events with a Student’s t likelihood and obtain a Gaussian approximation via a Laplace method [47, 85, 11, 65], while drawing on established Bayesian techniques for uncertainty estimation [85, 58, 33, 49, 18]. The resulting uncertainty-weighted fusion dynamically modulates each modality’s contribution, allowing high-resolution temporal cues from events to complement, rather than be suppressed by, the detailed spatial context of images.
Furthermore, our framework incorporates a sequential update mechanism and an iterative refinement process. The sequential update merges each new image–event observation with the prior fused state via inverse-variance (Kalman-gain) weights, closely mirroring the Kalman filter update step [64, 27]. Iterative refinement then targets the fine-grained residual errors that the fusion step cannot fully resolve—such as feature-level mismatches, modality-specific noise, and minor scale imbalances that arise when combining two streams with different noise profiles. By repeatedly estimating and subtracting these residuals, the refinement network progressively denoises and re-balances the fused representation, yielding a robust and cohesive latent state.
Our contributions can be summarized as follows:
  • Precision-Weighted Fusion Mechanism: Drawing inspiration from Bayesian inference and Kalman filtering, our framework employs inverse variance (precision) weighting to dynamically fuse image and event representations. This principled approach allows our model to prioritize the contribution of each modality based on its estimated uncertainty, addressing key limitations of prior fusion strategies that fail to adaptively account for heterogeneous uncertainty and often overly rely on dominant modalities.
  • Practical Integration of Synthetic Event Data: We present the successful use of synthetic event streams extracted from conventional videos, addressing the shortage of real-world event datasets. Our results demonstrate that this approach enables the integration of event modality into virtually any video-based dataset, substantially broadening the scope and applicability of multimodal anomaly detection.
  • Empirical Validation on Real-World Datasets: We validate our approach on four major real-world video anomaly detection datasets—UCF-Crime, XD-Violence, ShanghaiTech, and MSAD—achieving AUC scores of 88.67%, 87.63%, 97.98%, and 92.90%, respectively, surpassing state-of-the-art methods on each benchmark. Our results further demonstrate that integrating event modality enables the detection of motion-centric and transient anomalies that static-image-based approaches often fail to capture.
2 Related Work
Multimodal Data Fusion: Multimodal fusion aims to integrate complementary signals from diverse modalities such as vision, audio, and language [72, 87, 29]. Early approaches relied on simple feature concatenation [48], but recent advances have increasingly leveraged transformer-based architectures [62, 1, 24, 36] and contrastive pretraining [54, 23, 10] to capture complex cross-modal dependencies. Although attention-based fusion methods offer fine-grained alignment [38, 36], they often suffer from modality dominance, where one modality disproportionately influences the joint representation [72, 87]. In response, emerging works have explored using large language models (LLMs) to dynamically modulate modality usage based on context [56, 78, 25, 35, 13], thereby highlighting the need for adaptive and selective integration. Building on this direction, our work emphasizes uncertainty as an explicit control signal for fusion, enabling robust integration under modality imbalance and real-world noise conditions.
Utilization of Event-Based Data: Recent neuromorphic sensors and datasets [34, 50, 39] have renewed interest in the event modality, characterized by its asynchronous, sparse, and high-temporal-resolution nature that captures only salient changes [19, 79, 5, 4]. This distinctive property has been exploited across a wide range of tasks [57, 73, 44, 9, 52, 55], with recent methods leveraging CLIP [54] to achieve effective modality alignment between image and event data [71, 82, 31]. This alignment demonstrates the capability to process paired image and event inputs, thereby enabling the extraction of synthetic event data from videos that capture the motion cues essential for anomaly detection [31].
Video Anomaly Detection: The video anomaly detection literature has witnessed significant shifts towards weakly-supervised learning frameworks driven by multiple instance learning (MIL) strategies [59, 66, 80, 6, 46]. Recent transformer-based models, which predominantly rely on image-based representations, aim to detect subtle, transient anomalies in long video streams, yet emerging work suggests that incorporating richer temporal information can greatly improve performance [42, 21, 3]. In this regard, several recent studies have begun exploring the integration of additional modalities, including text inputs via large language models (LLMs), to provide extra contextual information that can aid temporal understanding [2, 13, 78, 74, 75, 45]. In the context of modality integration, the extraction of synthetic event data from video [31, 55, 84, 3] has emerged as a promising method for enriching temporal cues, focusing on capturing motion-centric signals rather than static background details. This refined focus enables models to better attend to the salient dynamics of the scene, thereby enhancing the detection of critical anomalous behavior.
Uncertainty Estimation & Bayesian Fusion: In parallel with advances in multimodal fusion, there has been a growing trend in incorporating uncertainty estimation directly into deep learning frameworks [33, 18, 51]. Recent works utilize Bayesian principles to quantify prediction uncertainty, thereby enabling models to weigh each modality’s contribution based on confidence measures [58, 47]. Approaches leveraging Monte Carlo dropout [18] and precise inverse variance weighting have been reported [58, 11], demonstrating improved performance in challenging conditions [51]. Our method adopts a Bayesian fusion framework where the inverse variance—computed via a Laplace approximation over a Student’s t noise model—is used to weight the fusion of image and event features [47, 65]. This not only aligns with recent trends but also provides a principled mechanism to enhance robustness, especially in the presence of heavy-tailed noise [65].
3 Uncertainty-Weighted Fusion for Multimodal Learning
Figure 1: Overview of IEF-VAD framework. Each video frame and its corresponding synthetic event representation are processed by CLIP encoders to obtain feature embeddings $z_{m}$. These are further encoded by modality-specific transformers $f_{m}$ to produce $\hat{z}_{m}$, which are then passed through projection heads $g_{m}$ and $h_{m}$ to estimate $\mu_{m}$ and $\sigma_{m}$. The estimated $\sigma_{m}$ is used to compute the uncertainty-aware fusion weight $w_{m}$, which is used to obtain the initial fused representation $\mu_{f}^{t,0}$. This is refined over $N$ iterative steps through a refinement network to produce the final output $\mu_{f}^{t,N}$.
In this work, we extract image and event embeddings ($z_{x},z_{e}$) separately from videos, assuming that both modalities observe the same underlying scene and share a common spatial structure while exhibiting complementary expressive and modality-specific features. $z_{x}$ aggregates complex attributes such as color, background, and spatial details, whereas $z_{e}$ encapsulates transient changes and temporal dynamics. To effectively combine these heterogeneous representations, we propose a fusion strategy that integrates Bayesian uncertainty estimation with Kalman filter update principles, where uncertainty quantifies the reliability of each value-level feature. Specifically, we estimate latent features and their associated uncertainties for each modality, and subsequently fuse them through a refinement network designed to iteratively correct residual noise, misalignments, and modality-specific artifacts, yielding a robust final representation $\mu_{f}$.
3.1 Modality Design and Uncertainty Estimation
We design each modality as a noisy observation influenced by both shared scene content and modality-specific factors.
$z_{m}=\mu_{m}+\delta_{m},\quad\delta_{m}\sim t_{\nu}(0,\Sigma_{m}),\quad m\in\{x,e\},$
where $\mu_{m}\!\in\!\mathbb{R}^{d}$ is the central estimate and $\Sigma_{m}\!=\!\operatorname{diag}(\sigma_{m}^{2})$ encodes value-level uncertainty. We choose the heavy-tailed Student’s-$t$ distribution because a purely Gaussian model underestimates uncertainty in the presence of outliers; its thicker tails make the fusion rule more conservative when the input is degraded (see Appendix A).
Given embedding sequences $z_{m}\!\in\!\mathbb{R}^{B\times T\times D}$, we transpose them to $(T,B,D)$, pass each through a modality-specific transformer $f_{m}$, and transpose back, yielding $\hat{z}_{m}=(f_{m}(z_{m}^{\top}))^{\top}$.
Layer normalization aligns scales across modalities. This ensures that the downstream uncertainty estimation operates on comparable feature scales, improving numerical stability when computing variances. Linear projection heads $g_{m},h_{m}$ then predict the posterior mean and log-variance:
$\mu_{m}=g_{m}(\hat{z}_{m}),\quad\log\sigma_{m}^{2}=h_{m}(\hat{z}_{m}).$ (1)
Predicting $\log\sigma_{m}^{2}$ guarantees positivity and numerical stability. Conceptually, $z_{m}$ is sampled from the posterior $t_{\nu}(\mu_{m},\Sigma_{m})$, where $\mu_{m}$ represents the fused central tendency and $\sigma_{m}^{2}$ quantifies epistemic uncertainty. These uncertainty estimates drive the Kalman-style update and the subsequent refinement loop, enabling dynamic, reliability-aware fusion that adapts to modality quality in real time.
3.2 Inverse Variance Calculation
In this stage, our goal is to quantify the confidence in each modality’s predictions by computing the inverse variance. Recall that our model predicts the log-variance values (Eq. 1), which are then exponentiated to recover the variance. In the case of a Gaussian noise model, this is given by $\sigma_{m}^{2}=\exp\big{(}\log\sigma_{m}^{2}\big{)}$. For the Student’s t noise model, we first apply the Laplace approximation to obtain the effective variance (see Appendix B). The Student’s t probability density function (up to a normalization constant) is given by Eq. 2, where $\sigma^{2}$ is the variance parameter for the underlying Gaussian scale, and $\nu$ is the degree of freedom.
$p(\delta)\propto\left(1+\frac{\delta^{2}}{\nu\sigma^{2}}\right)^{-\frac{\nu+1}{2}},\quad\log p(\delta)=-\frac{\nu+1}{2}\log\left(1+\frac{\delta^{2}}{\nu\sigma^{2}}\right)+C.$ (2)
Since the mode of the distribution occurs at $\delta=0$, we perform a second-order Taylor expansion of the logarithm around $\delta=0$. For small $x$, recall that $\log(1+x)\approx x$ for $|x|\ll 1$. In our case, set $x=\delta^{2}/(\nu\sigma^{2})$. Then, for small $\delta$, and log-density we have
$\log\left(1+\frac{\delta^{2}}{\nu\sigma^{2}}\right)\approx\frac{\delta^{2}}{\nu\sigma^{2}},\quad\log p(\delta)\approx-\frac{\nu+1}{2}\frac{\delta^{2}}{\nu\sigma^{2}}+C=-\frac{\nu+1}{2\nu\sigma^{2}}\delta^{2}+C.$
The Laplace approximation approximates a probability density near its mode by a Gaussian distribution. The log-density of a Gaussian with variance $\tilde{\sigma}^{2}$ is given by $\log p_{G}(\delta)=-\frac{1}{2\tilde{\sigma}^{2}}\delta^{2}+C.$ By matching the quadratic terms in the Taylor expansion, we set $\frac{1}{2\tilde{\sigma}^{2}}=\frac{\nu+1}{2\nu\sigma^{2}}$. This immediately implies $\tilde{\sigma}^{2}=\frac{\nu}{\nu+1}\sigma^{2}.$ Taking the logarithm of both sides gives
$\log\tilde{\sigma}^{2}=\log\sigma^{2}+\log\left(\frac{\nu}{\nu+1}\right).$ (3)
This derivation shows that, under the Laplace approximation, the effective variance used in downstream computations is scaled by the factor $\frac{\nu}{\nu+1}$, reflecting the heavy-tailed nature of the Student’s t distribution. This effective variance is then used in place of the original variance $\sigma^{2}$ when computing inverse variance weights and other related measures, ensuring that the fusion process properly accounts for the increased uncertainty due to heavy-tailed noise.
The variance (or effective variance in the Student’s t case) represents the uncertainty associated with the prediction; lower values indicate higher confidence. To leverage this notion of confidence in a manner consistent with Bayesian principles, we compute the inverse variance as a measure of precision. Specifically, for the Student’s t model, we use the effective variance:
$w_{m}=\frac{1}{\tilde{\sigma}_{m}^{2}+\epsilon}.$ (4)
where $\epsilon$ is a small positive constant added for numerical stability, ensuring that we do not encounter division by zero. These weights, $w_{x}$ and $w_{e}$, essentially serve as confidence scores—modalities with lower uncertainty (i.e., lower $\sigma^{2}$ or $\tilde{\sigma}^{2}$) yield higher weights and, consequently, contribute more significantly in subsequent fusion steps. This approach is theoretically grounded in Bayesian inference (detailed in Appendix C), where the inverse variance (or precision) is used to weight the contributions of different measurements according to their reliability. As a result, by explicitly modeling and incorporating the inverse variance—or the effective inverse variance under the Student’s t assumption—the fusion process becomes more robust, effectively balancing the contributions of each modality based on their estimated uncertainties.
3.3 Uncertainty-Weighted Fusion
We compute the fused representation $\mu_{f}\in\mathbb{R}^{B\times T\times D}$ by taking a weighted average of the modality-specific means according to their computed confidence scores Eq. 4.
$\mu_{f}=\frac{w_{x}\mu_{x}+w_{e}\mu_{e}}{w_{x}+w_{e}}$
In the case of the Student’s t distribution, the inverse variance weights are computed based on the effective variances obtained via a Laplace approximation around the mode of the distribution. Specifically, by applying the logarithm to the effective variance as derived in Eq. 3, and then exponentiating, we have
$\tilde{\sigma}_{m}^{2}=\exp\left(\log\sigma_{m}^{2}+\log\left(\frac{\nu}{\nu+1}\right)\right)$
Thus, in both cases, modalities with lower uncertainty (i.e., lower variance or effective variance) yield higher precision scores and contribute more significantly in the fusion process.
The rationale behind this formulation is twofold. First, by weighting each modality’s mean by its (effective) inverse variance, we ensure that modalities with higher confidence have a greater impact on the final fused representation. This is a direct application of Bayesian fusion principles, where the posterior estimate of a latent variable is a precision-weighted average of the individual estimates. Second, this fusion strategy closely mirrors the update step in the Kalman filter—a well-established method for sequential data fusion—where the Kalman gain, derived from the inverse variances, dictates the contribution of each measurement in updating the state estimate, as further detailed in Appendix C.
3.4 Time Step Sequential Update
Building upon our uncertainty-weighted fusion framework (Figure 1), we extend the method to handle the time steps ($T$) by incorporating temporal dependencies in a sequential update process. The intuition is analogous to the recursive estimation in Kalman filtering, except that here we account for the heavy-tailed nature of the Student’s t noise via a Laplace approximation.
Under the Student’s t model, the predicted variance is corrected to obtain an effective variance. Specifically, if the predicted log-variance is $\log\sigma^{2}_{t}$ at time $t$, the effective variance and state variance at the previous time step are given by
$\tilde{\sigma}_{t}^{2}=\exp\left(\log\sigma_{t}^{2}+\log\left(\frac{\nu}{\nu+1}\right)\right)$ (5)
At the initial time step ($t=0$), we set the state and its effective uncertainty directly from the first input: $\mu_{f}^{0}=\mu^{0},\quad\tilde{\sigma}_{f,0}^{2}=\tilde{\sigma}_{0}^{2}$. For each subsequent time step ($t\geq 1$), we update the state estimate by fusing the previous state with the current input. To do so, we compute the inverse effective variance weights. The weight for the previous state is given by $w_{f}^{t-1}=1/(\tilde{\sigma}_{f,t-1}^{2}+\epsilon)$ and the weight for the current input is $w^{t}=1/(\tilde{\sigma}_{t}^{2}+\epsilon)$, where $\epsilon$ is a small constant for numerical stability. The updated state and its effective uncertainty are computed as:
$\mu_{f}^{t}=\frac{w_{f}^{t-1}\mu_{f}^{t-1}+w^{t}\mu^{t}}{w_{f}^{t-1}+w^{t}},\quad\tilde{\sigma}_{f,t}^{2}=\frac{1}{w_{f}^{t-1}+w^{t}},\quad t\geq 1.$
This sequential update process naturally extends the static fusion methodology by incorporating temporal continuity, using effective variances that account for the heavy-tailed nature of the Student’s t noise. Just as in the static case—where each modality is weighted according to its (effective) precision—the sequential update fuses the previous state with new observations based on their relative effective uncertainties. This approach enhances the robustness and consistency of the state estimates over time, effectively capturing both current observations and historical context.
3.5 Iterative Refinement of Fused State
After performing sequential fusion based on uncertainty-weighted averaging of multimodal latent representations, residual noise and subtle discrepancies between modalities inevitably persist in the fused latent state. Specifically, since each modality introduces distinct noise profiles and uncertainty estimates, their combination inherently results in microscopic residual errors. To systematically address these fusion-induced residuals, we introduce an iterative refinement procedure, inspired by iterative denoising methods such as the Denoising Diffusion Probabilistic Model [28].
Formally, starting from an initial fused state $\mu_{f}^{0}$ obtained after the sequential uncertainty-weighted update, we iteratively predict and subtract the residual error. At each refinement iteration $r$ ($r=0,1,\dots,N$), a dedicated refinement network $F(\cdot)$ estimates the residual error $\Delta\mu_{f}^{t,r}=F(\mu_{f}^{t,r})$ based solely on the current fused state. The estimated residual represents the remaining discrepancy between the current fused representation and the underlying true latent state. To ensure stable and progressive refinement without overshooting, the state is updated using a fixed attenuation parameter $\lambda_{r}\in(0,1)$:
$\mu_{f}^{t,r+1}=\mu_{f}^{t,r}-\lambda_{r}\Delta\mu_{f}^{t,r}.$
This attenuation parameter modulates the magnitude of each residual correction, preventing over-correction and ensuring stable convergence. As iterative refinement progresses, residual errors decrease monotonically, gradually aligning the fused latent representation closer to the underlying true latent posterior distribution. Thus, iterative refinement serves as a theoretically and empirically justified mechanism to denoise and enhance the multimodal fused representation, yielding a more precise and reliable latent embedding suitable for subsequent downstream tasks (detailed in Appendix D).
3.6 Video Anomaly Detection with Loss Functions
Classification Loss ($\mathcal{L}_{\text{cls}}$): After obtaining the final refined fused representation $\mu_{f}^{t,N}\in\mathbb{R}^{B\times T\times D}$, we perform binary classification at each time step to detect anomalies. Specifically, we apply a lightweight classification head ($H$) to $\mu_{f}^{t,N}$ to produce logits:
$\hat{y}=H(\mu_{f}^{t,N})\in\mathbb{R}^{B\times T\times 1}.$
The logits are then passed through a sigmoid activation to obtain anomaly probabilities and a binary cross-entropy loss $\mathcal{L}_{\text{cls}}$ is computed against the ground-truth labels.
$\mathcal{L}_{\text{cls}}=-\frac{1}{BT}\sum_{b=1}^{B}\sum_{t=1}^{T}\left[y_{b,t}\log(\hat{y}_{b,t})+(1-y_{b,t})\log(1-\hat{y}_{b,t})\right],$
To accommodate variability in temporal granularity and enable weakly supervised learning without requiring precise frame-level annotations, we adopt a segment-based aggregation strategy following [59]. Specifically, the temporal sequence is divided into non-overlapping segments of 16 frames. Let length denote the number of valid time steps. Then, the number of segments $k$ is computed as $k=\left\lfloor\frac{\text{length}}{16}\right\rfloor+1$, and the predictions within each segment are averaged to yield an instance-level prediction [30, 59].
KL Divergence Loss ($\mathcal{L}_{\text{KL}}$): Under the Student’s t noise model, each modality’s measurement noise is assumed to follow $\delta_{m}\sim t_{\nu}(0,\sigma_{m}^{2})$, where $\nu$ denotes the degrees of freedom. Direct computation of the KL divergence between a Student’s t distribution and the standard normal prior is intractable due to the heavy-tailed nature of the distribution. To address this, we leverage the effective variance $\tilde{\sigma}_{m}^{2}$ Eq. 5 obtained via Laplace approximation (Eq. 3) and approximate the latent distribution as a Gaussian $\mathcal{N}(\mu_{m},\tilde{\sigma}_{m}^{2})$. The resulting closed-form KL divergence with respect to $\mathcal{N}(0,I)$ is given by (see Appendix E for detailed derivation):
$\mathcal{L}_{\text{KL}}=KL\big{(}\mathcal{N}(\mu_{m},\tilde{\sigma}_{m}^{2})\|\mathcal{N}(0,I)\big{)}=\frac{1}{2}\big{(}\tilde{\sigma}_{m}^{2}+\mu_{m}^{2}-1-\log\tilde{\sigma}_{m}^{2}\big{)}$
This regularization encourages the latent distribution to remain close to the prior, providing robustness even under heavy-tailed noise conditions.
Regularization Loss ($\mathcal{L}_{\text{reg}}$): To ensure that the latent representations from both modalities are aligned both in direction and magnitude, we introduce a regularization loss. This term comprises two components: one that minimizes the angular difference between $\mu_{x}$ and $\mu_{e}$ (using cosine similarity), and another that penalizes discrepancies in their norms. Formally, it is defined as
$\mathcal{L}_{\text{reg}}=\lambda_{1}(1-\cos(\mu_{x},\mu_{e}))+\lambda_{2}|\|\mu_{x}\|-\|\mu_{e}\||$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that balance the contributions of the two terms.
$\mathcal{L}_{\text{reg}}$ promotes a shared latent space across modalities by jointly regularizing the direction and magnitude of their embeddings. The cosine term enforces semantic alignment by minimizing angular deviations, while the norm consistency term balances embedding scales to prevent modality domination. Together, these constraints enable stable and coherent multimodal fusion.
Overall Loss: The final loss function is a sum of the classifier loss, the KL divergence losses for both modalities, and the regularization loss:
$\mathcal{L}=\mathcal{L}_{\text{cls}}+\sum_{m}\mathcal{L}_{\text{KL}}^{m}+\mathcal{L}_{\text{reg}}$ (6)
This composite loss ensures that the model not only produces accurate binary predictions at each time step but also learns latent representations that are both robust and well-regularized. The use of the effective variance derived via Laplace’s approximation allows us to maintain a closed-form KL divergence expression, thereby combining the robustness of the Student’s t model with the computational efficiency of Gaussian-based methods. An ablation study on the loss components is provided in Appendix G.
4 Experiment Results
We evaluate our method on four public video anomaly detection datasets—UCF-Crime [59], XD-Violence [66], ShanghaiTech [40], and MSAD [86]—covering diverse real-world surveillance scenarios. Following standardized preprocessing and independent transformer encoding for each modality (see Appendix F), we perform evaluations using AUC and AP metrics. Ablation studies presented in Appendix G analyze the sensitivity of performance to key hyperparameters ($\nu$, $N$, $\lambda_{\text{ref}}$, and $\epsilon$) and investigate the effects of loss configuration choices (Eq. 6). Robustness to outlier injection and the behavior of uncertainty weights under perturbations are evaluated, with additional uncertainty-aware metrics, including KL divergence and Brier score, reported in Appendix H. Our method consistently demonstrates strong performance by effectively leveraging complementary spatial and temporal cues from image and event modalities in a weakly supervised setting.
4.1 Real-World Anomaly Detection in Surveillance Video
Method UCF-Crime [59] XD-Violence [66] Shanghai-Tech [40] Method MSAD [86]
AUC (%) Ano-AUC (%) AP (%) AUC (%) AUC (%)
Sultani et al. [59] 84.14 63.29 75.18 91.72 MIST (I3D) [15] 86.65
Wu et al. [66] 84.57 62.21 80.00 95.24 MIST (SwinT) [15] 85.67
AVVD [67] 82.45 60.27 - - UR-DMU [81] 85.02
RTFM [61] 85.66 63.86 78.27 97.21 UR-DMU (SwinT) [81] 72.36
UR-DMU [81] 86.97 68.62 81.66 97.57 MGFN (I3D) [8] 84.96
UMIL [46] 86.75 68.68 - 96.78 MGFN (SwinT) [8] 78.94
VadCLIP [70] 88.02 70.23 84.51 97.49 TEVAD (I3D) [7] 86.82
STPrompt [69] 88.08 - - 97.81 TEVAD (SwinT) [7] 83.6
OVVAD [68] 86.40 - 66.53 96.98 EGO [12] 87.36
IEF-VAD (Gaussian) 88.11 $\pm$ 0.28 70.48 $\pm$ 0.66 87.18 $\pm$ 0.55 97.91 $\pm$ 0.08 IEF-VAD (Gaussian) 92.27 $\pm$ 0.34
IEF-VAD (Student-T) 88.67 $\pm$ 0.45 71.50 $\pm$ 1.02 87.63 $\pm$ 0.54 97.98 $\pm$ 0.07 IEF-VAD (Student-T) 92.90 $\pm$ 0.27
Table 1: Comparison of various methods on multiple anomaly detection benchmarks, including UCF-Crime [59], XD-Violence [66], ShanghaiTech [40], and MSAD [86]. All metrics are reported as the mean (±1 standard deviation) of 10 runs. Our approach (IEF-VAD) consistently achieves the highest performance across datasets, demonstrating its robustness for video anomaly detection.
Table 1 shows that IEF-VAD outperforms all prior weakly supervised detectors on every benchmark. On UCF-Crime [59], the Gaussian variant already matches the best published AUC (88.11%), while the Student’s t extension lifts AUC to 88.67% and raises the anomaly-focused AUC (Ano-AUC) to 71.50%—a $\approx$ 1.3 pp absolute gain over the previous record (70.23% of VadCLIP [70]). Similar trends appear on XD-Violence [66] (87.63 AP) and ShanghaiTech [40] (97.98 AUC), where both variants surpass the strongest competitors. On the more recent MSAD [86] benchmark, our Student’s-t model achieves 92.90 AUC, exceeding the best I3D-based baseline (86.82 AUC) by over 6 pp. Standard-deviation margins indicate that gains are statistically consistent across 10 independent runs.
These results confirm two key insights of our framework. First, value-level uncertainty weighting enables practical exploitation of the heterogeneous synthetic-event modality, turning its complementary cues into measurable performance gains whenever the RGB stream is degraded. Second, modelling heavy-tailed noise via a Student’s-t likelihood yields further, systematic improvements, especially on long-tailed datasets (UCF-Crime [59], XD-Violence [66]) where RGB frames frequently contain motion blur or illumination changes. The consistent gains across four diverse datasets—with no task-specific tuning—underscore the generality of uncertainty-aware fusion and highlight its promise for real-world surveillance settings in which sensor quality and scene dynamics vary unpredictably.
Figure 2: Radar charts showing per-class anomaly detection performance (AUC and AP) for image-only (blue), event-only (orange), and fused (green) approaches. Each radial axis represents an anomaly category, and values are normalized per class by the maximum score. The fused approach consistently covers a larger area, highlighting improved detection across anomaly types.
These trends are further illustrated in Figure 2, which provides a class-wise comparison of image-only, event-only, and fused approaches. In most anomaly classes, the image-based modality achieves higher performance than the event-based modality, reflecting the substantial differences in the underlying information each modality encodes. However, in anomaly categories that align more closely with the characteristics of $z_{e}$ (e.g., Fighting, Assault in UCF-Crime [59]), the event-based modality outperforms the image-based one, confirming our initial intuition (Figure 2). By fusing $z_{x}$ and $z_{e}$, our proposed method consistently achieves higher detection accuracy than either single modality alone. In particular, this fusion harnesses the strengths of image while absorbing the advantages of event for those classes where it excels, leading to further performance gains in categories already well-handled by image. Consequently, as illustrated in Figure 2, the fused approach covers a broader area in the radial plots, surpassing the capacity of single-modality baselines in most anomaly classes (see Appendix F for detailed numbers).
4.2 Fusion: Uncertainty Behavior Under Value-Level Masking
Figure 3: Change in image-side uncertainty weights $\Delta w_{x}$ under modality-specific masking perturbations across four datasets. The horizontal axis $i$ denotes the latent dimension index, and the vertical axis $\rho$ indicates the proportion of values in $z_{x}$ or $z_{e}$ that are masked to zero. Each surface visualizes $\Delta w_{x}(i,\rho)=w_{x}^{\text{masked}}(i)-w_{x}^{\text{clean}}(i)$ for a given masking ratio. The top row corresponds to masking applied to $z_{x}$ (image modality), while the bottom row applies masking to $z_{e}$ (event modality), with both measuring the resulting change in $w_{x}$. Positive values (blue) indicate increased confidence in the image modality under corruption, while negative values (red) reflect a reduction. The non-uniform patterns across $i$ highlight dimension-specific responses to value-level degradation.
To investigate how the uncertainty weights $w_{m}$ respond to modality-specific degradation, we perturb a single modality by masking a random subset of its latent features $z_{m}\in\mathbb{R}^{B\times T\times D}$. A masked sequence is defined as
$z_{m,i}^{\text{masked}}=\begin{cases}0,&i\in\mathcal{I}_{\rho}\\[4.0pt] z_{m,i},&\text{otherwise}\end{cases},\qquad|\mathcal{I}_{\rho}|=\rho D.$
where $\rho\in(0,1)$ denotes the masking ratio and $i$ indexes feature dimensions. We then measure the change in the image-side uncertainty weight, $\Delta w_{x}=w_{x}^{\text{masked}}-w_{x}^{\text{clean}}$, with the normalisation $w_{x}+w_{e}=1$ enforced. We prefer masking over additive noise because it produces a deterministic, localised degradation that preserves the Student-$t$ noise assumption and avoids the non-linear propagation artifacts that Gaussian perturbations introduce in attention layers, yielding a cleaner causal probe of the learned uncertainty mechanism.
Figure 3 plots $\Delta w_{x}$ for three masking ratios ($\rho\in{0.05,0.10,0.20}$) on four datasets. The top row masks $z_{x}$, consistently decreasing $w_{x}$—indicating higher image uncertainty—whereas the bottom row masks $z_{e}$, symmetrically increasing $w_{x}$. The magnitude of $|\Delta w_{x}|$ grows with $\rho$, showing that the fusion network scales its confidence shift with corruption severity.
Value-level curves expose pronounced heterogeneity: some dimensions (e.g., indices 12–27) react strongly, whereas others remain flat. We attribute this to distributed encoding: dimensions dominated by stable appearance cues are robust, while those capturing transient, modality-specific dynamics are fragile. The non-uniform yet directionally consistent responses provide empirical evidence that the model estimates uncertainty on a fine-grained basis rather than collapsing it into a scalar. Supplementary statistics, including KL divergence and Brier score, are reported in Appendix H.
5 Conclusion
IEF-VAD proves that motion-centric synthetic event streams distilled from ordinary videos can be fused with RGB features to push video-anomaly detection beyond the limits of frame-based models. The approach establishes new state-of-the-art scores on UCF-Crime, XD-Violence, ShanghaiTech, and MSAD, while a value-level masking study shows its precision weights shift adaptively across latent dimensions—making the fusion logic transparent and reliable. We expect these findings to catalyse broader use of synthetic event data and to drive future work on uncertainty-aware multimodal learning for real-time surveillance.
Limitations and Future Work. While our study advances uncertainty-guided fusion, several limitations invite further investigation. First, approximating each modality’s noise with a diagonal covariance neglects cross-feature correlations; future work could adopt lightweight structured or low-rank covariance estimators to enrich value-level fusion. Second, the fixed regularization weights $\lambda_{1,2}$ and Student-t degrees-of-freedom $\nu$ limit adaptability; jointly learning or meta-learning these parameters would enable the model to recalibrate trust as modality quality shifts. Third, replacing simple weight averaging with uncertainty-conditioned cross-modal attention, hierarchical gating, or iterative message passing may yield a more expressive fusion mechanism that leverages inter-modal disagreement as informative signal, paving the way for stronger anomaly detection and broader multimodal perception capabilities.
Acknowledgments
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, and by the Semiconductor Research Corporation (SRC). Additional support was provided by the Office of Naval Research through the Young Investigator Program Award and Grants #N00014-21-1-2225 and #N00014-22-1-2067, as well as the Army Research Office under Grant #W911NF2410360. This research was also supported by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, and through generous gifts from Xilinx and Cisco.
Appendix A Bounded Influence of Student’s t-Noise
Proposition 1 (Robustness of Student’s t to Outliers).
Let $\delta$ be a noise or residual term drawn from a univariate Student’s t distribution with $\nu>0$ degrees of freedom, location 00, and scale $\sigma$:
$\delta\sim t_{\nu}(0,\sigma^{2}).$
The probability density function of $\delta$ is proportional to
$p(\delta)\;\propto\;\Bigl{(}1+\tfrac{\delta^{2}}{\nu\,\sigma^{2}}\Bigr{)}^{-\tfrac{\nu+1}{2}}.$
Hence, the negative log-likelihood (omitting constant terms that do not depend on $\delta$) is
[…]
Differentiating with respect to $\delta$ yields the score function
[…]
As $\lvert\delta\rvert\to\infty$, the denominator $\nu\,\sigma^{2}+\delta^{2}$ is dominated by $\delta^{2}$, so
[…]
Thus, the derivative (i.e., the slope or ‘pull’ of the residual on the log-likelihood) remains bounded and actually tends to zero for large outliers.
In contrast, consider a Gaussian noise model, $\delta\sim\mathcal{N}(0,\sigma^{2})$. The corresponding negative log-likelihood is
[…]
whose derivative is
$\frac{d}{d\delta}\bigl{[}-\log p(\delta)\bigr{]}\;=\;\frac{\delta}{\sigma^{2}},$
which grows unboundedly as $\lvert\delta\rvert\to\infty$. Therefore, large outliers in a Gaussian model have much stronger influence on parameter estimation, making it less robust to extreme residuals.
Hence, the bounded slope in the Student-$t$ model demonstrates greater robustness against outliers: as $\lvert\delta\rvert$ becomes large, its influence on parameter updates (through gradient-based or maximum likelihood methods) remains finite. This property is central to why Student-$t$-based methods are often preferred in situations where occasional extreme values are expected.
Appendix B Laplace Approximation for the Student’s t Noise Model
We derive an effective variance for the Student’s t noise model via the Laplace approximation. Our goal is to approximate the heavy-tailed log-density by a quadratic (Gaussian) form in the vicinity of its mode, thereby obtaining an effective variance that scales the underlying Gaussian variance by a factor of $\nu/(\nu+1)$.
Student’s t Distribution
The probability density function of the Student’s t distribution with degrees of freedom $\nu$, location parameter $\mu$, and scale parameter $s>0$ is given by
$f(x;\nu,\mu,s)=\frac{\Gamma\!\left(\frac{\nu+1}{2}\right)}{s\,\sqrt{\nu\pi}\,\Gamma\!\left(\frac{\nu}{2}\right)}\left(1+\frac{1}{\nu}\left(\frac{x-\mu}{s}\right)^{2}\right)^{-\frac{\nu+1}{2}},$
where:
  • $\Gamma(\cdot)$ denotes the Gamma function.
  • $\nu>0$ is the degrees of freedom, controlling the heaviness of the tails.
  • $\mu\in\mathbb{R}$ is the location parameter (here, assumed to be zero in our derivation).
  • $s>0$ is the scale parameter; setting $\sigma^{2}=s^{2}$ allows us to interpret $\sigma^{2}$ as the variance of the underlying Gaussian scale.
In our derivation we assume $\mu=0$ for simplicity. Ignoring the normalization constant, the unnormalized density can then be written as
$p(\delta)\propto\left(1+\frac{\delta^{2}}{\nu\,\sigma^{2}}\right)^{-\frac{\nu+1}{2}},$
where $\delta\in\mathbb{R}$ represents the noise term.
Derivation via Laplace Approximation
We now present a detailed derivation of the effective variance via the Laplace approximation.
Proposition 2 (Effective Variance under the Laplace Approximation).
Let $\delta\in\mathbb{R}$ be a noise term with distribution (ignoring the normalization constant)
$p(\delta)\propto\left(1+\frac{\delta^{2}}{\nu\,\sigma^{2}}\right)^{-\frac{\nu+1}{2}},$
where $\sigma^{2}$ is the scale (variance) parameter of the underlying Gaussian and $\nu>0$ is the degrees of freedom. Then, by performing a second-order Taylor expansion of the log-density about its mode at $\delta=0$, the local approximation is equivalent to that of a Gaussian distribution with effective variance
$\tilde{\sigma}^{2}=\frac{\nu}{\nu+1}\,\sigma^{2}.$
We start with the unnormalized density and take the natural logarithm to obtain the log-density as
$p(\delta)\propto\left(1+\frac{\delta^{2}}{\nu\,\sigma^{2}}\right)^{-\frac{\nu+1}{2}},\quad\log p(\delta)=-\frac{\nu+1}{2}\,\log\!\left(1+\frac{\delta^{2}}{\nu\,\sigma^{2}}\right)+C,$
where $C$ is a constant independent of $\delta$.
Since the mode of $p(\delta)$ occurs at $\delta=0$, we perform a Taylor expansion of $\log p(\delta)$ around $\delta=0$. For small $x$, we have the approximation
[…]
Thus, for small $\delta$,
$\log\!\left(1+\frac{\delta^{2}}{\nu\,\sigma^{2}}\right)\approx\frac{\delta^{2}}{\nu\,\sigma^{2}}.$
Substituting this into the log-density expression yields
$\log p(\delta)\approx-\frac{\nu+1}{2}\,\frac{\delta^{2}}{\nu\,\sigma^{2}}+C=-\frac{\nu+1}{2\nu\,\sigma^{2}}\,\delta^{2}+C.$
Now, consider the log-density of a Gaussian distribution with mean zero and variance $\tilde{\sigma}^{2}$:
$\log p_{G}(\delta)=-\frac{1}{2\tilde{\sigma}^{2}}\,\delta^{2}+C^{\prime},$
where $C^{\prime}$ is a constant independent of $\delta$. To match the local quadratic approximation of $\log p(\delta)$, we equate the coefficients of $\delta^{2}$:
$\frac{1}{2\tilde{\sigma}^{2}}=\frac{\nu+1}{2\nu\,\sigma^{2}}.$
Multiplying both sides by 2 gives
$\frac{1}{\tilde{\sigma}^{2}}=\frac{\nu+1}{\nu\,\sigma^{2}}.$
Taking reciprocals, we obtain
$\tilde{\sigma}^{2}=\frac{\nu}{\nu+1}\,\sigma^{2}.$
Appendix C Theoretical Foundations of Inverse Variance Weighting in Bayesian Inference
In this appendix, we provide a detailed derivation and theoretical justification of the inverse variance (or precision) weighting scheme used in our fusion model. This method is firmly rooted in Bayesian inference, where each measurement contributes to the estimation of the latent variable based on its reliability.
C.1 Measurement Fusion Under Gaussian Noise
Assume that we wish to estimate a latent variable $z$ from two independent noisy measurements $z_{x}$ and $z_{e}$. Each measurement is modeled as:
$z_{m}=\mu_{m}+\delta_{m},\quad\delta_{m}\sim\mathcal{N}(0,\sigma_{m}^{2}),\quad m\in\{x,e\}.$
Here, $\mu_{m}$ represents the central estimate predicted by modality $m$, and $\sigma_{m}^{2}$ is the associated uncertainty (variance). The likelihood of observing $z_{m}$ given $z$ is then
$p(z_{m}\mid z)\propto\exp\!\left(-\frac{(z_{m}-z)^{2}}{2\sigma_{m}^{2}}\right).$
Assuming a flat prior $p(z)$, the posterior is proportional to the product of the likelihoods:
$p(z\mid z_{x},z_{e})\propto p(z_{x}\mid z)\,p(z_{e}\mid z).$
Taking the logarithm, we obtain the joint log-likelihood:
$\log p(z\mid z_{x},z_{e})=-\frac{(z_{x}-z)^{2}}{2\sigma_{x}^{2}}-\frac{(z_{e}-z)^{2}}{2\sigma_{e}^{2}}+C.$
Differentiating with respect to $z$ to find the maximum a posteriori (MAP) estimate $\hat{z}$:
$\frac{\partial}{\partial z}\left[-\frac{(z_{x}-z)^{2}}{2\sigma_{x}^{2}}-\frac{(z_{e}-z)^{2}}{2\sigma_{e}^{2}}\right]=0,\quad\Rightarrow\quad\frac{z_{x}-z}{\sigma_{x}^{2}}+\frac{z_{e}-z}{\sigma_{e}^{2}}=0.$
Rearranging terms gives:
$z\left(\frac{1}{\sigma_{x}^{2}}+\frac{1}{\sigma_{e}^{2}}\right)=\frac{z_{x}}{\sigma_{x}^{2}}+\frac{z_{e}}{\sigma_{e}^{2}},$
and hence the fused estimate is:
$\hat{z}=\frac{\frac{z_{x}}{\sigma_{x}^{2}}+\frac{z_{e}}{\sigma_{e}^{2}}}{\frac{1}{\sigma_{x}^{2}}+\frac{1}{\sigma_{e}^{2}}}.$
This derivation shows that the optimal fusion under Gaussian noise is achieved by weighting each measurement with its inverse variance $w_{m}=\frac{1}{\sigma_{m}^{2}}$.
C.2 Bayesian Justification via the Kalman Filtering Framework
The inverse variance weighting rule is further supported by the Bayesian update formulations seen in Kalman filtering. Consider a scenario where a prior estimate $\hat{z}^{-}$ with variance $\sigma^{-}$ is updated with a measurement $z_{m}$ having uncertainty $\sigma_{m}^{2}$. The Kalman update is given by:
$\hat{z}=\hat{z}^{-}+K(z_{m}-\hat{z}^{-}),$
where the Kalman gain $K$ is:
$K=\frac{\sigma^{-}}{\sigma^{-}+\sigma_{m}^{2}}.$
A measurement with lower uncertainty (higher precision) results in a larger Kalman gain, thus exerting a greater influence on the updated state. Extending this idea to the fusion of multiple modalities, the final fused estimate can be expressed as:
$\mu_{f}=\frac{w_{x}\mu_{x}+w_{e}\mu_{e}}{w_{x}+w_{e}},$
which is exactly the precision-weighted average obtained via the MAP estimation under the assumed likelihood models.
C.3 Fusion Formula: Dual Theoretical Foundations
Our fusion formula is derived based on two theoretical foundations: Bayesian inference and the Kalman filter. First, assume that the two modalities provide independent estimates of the same latent variable $z$. For the Gaussian case, the estimates are given by
$p(z\mid\mu_{i},\sigma_{i}^{2})=\mathcal{N}(z;\mu_{i},\sigma_{i}^{2})\quad\text{and}\quad p(z\mid\mu_{e},\sigma_{e}^{2})=\mathcal{N}(z;\mu_{e},\sigma_{e}^{2}).$
Because these estimates are independent, the joint likelihood (or the unnormalized posterior under a uniform prior) is proportional to the product of the two Gaussians:
$p(z\mid\mu_{i},\mu_{e})\propto\exp\left(-\frac{||z-\mu_{i}||^{2}}{2\sigma_{i}^{2}}\right)\cdot\exp\left(-\frac{||z-\mu_{e}||^{2}}{2\sigma_{e}^{2}}\right).$
By combining the exponents and completing the square, we find that the value of $z$ that maximizes the posterior—i.e., the fused mean—is given by
$\mu_{f}=\frac{\mu_{i}/\sigma_{i}^{2}+\mu_{e}/\sigma_{e}^{2}}{1/\sigma_{i}^{2}+1/\sigma_{e}^{2}}.$
For the Student’s $t$ model, we replace $\sigma_{i}^{2}$ and $\sigma_{e}^{2}$ with their effective counterparts $\tilde{\sigma_{i}}^{2}$ and $\tilde{\sigma_{e}}^{2}$, as derived via the Laplace approximation (see below), leading to the same formulation in terms of the inverse variances.
From the perspective of the Kalman filter, suppose that one modality provides a prediction $\mu_{i}$ with variance $\sigma_{i}^{2}$ (or $\tilde{\sigma_{i}}^{2}$ in the Student’s $t$ case) and another modality provides a prediction $\mu_{e}$ with variance $\sigma_{e}^{2}$ (or $\tilde{\sigma_{e}}^{2}$ in the Student’s $t$ case). The Kalman filter update for the state estimates is given by
$\mu_{f}=\mu_{i}+K(\mu_{e}-\mu_{i}),$
where the Kalman gain $K$ is defined as
$K=\frac{\tilde{\sigma_{i}}^{2}}{\tilde{\sigma_{i}}^{2}+\tilde{\sigma_{e}}^{2}},$
which leads to an equivalent expression for the fused mean:
$\mu_{f}=\frac{\tilde{\sigma_{e}}^{2}\mu_{i}+\tilde{\sigma_{i}}^{2}\mu_{e}}{\tilde{\sigma_{i}}^{2}+\tilde{\sigma_{e}}^{2}}.$
By defining the inverse variance weights as $w_{i}=1/(\sigma_{i}^{2}+\epsilon)$ (or $w_{i}=1/(\tilde{\sigma_{i}}^{2}+\epsilon)$ for the Student’s $t$) and $w_{e}=1/(\sigma_{e}^{2}+\epsilon)$ (or $w_{e}=1/(\tilde{\sigma_{e}}^{2}+\epsilon)$ for the Student’s $t$), our fusion formula becomes
$\mu_{f}=\frac{w_{i}\mu_{i}+w_{e}\mu_{e}}{w_{i}+w_{e}}.$
Thus, both the Bayesian derivation and the Kalman filter interpretation lead to the same uncertainty-weighted fusion formula, with the only difference being that for the Student’s $t$ noise model we use a corrected (effective) variance.
To summarize, our fusion formula is derived based on two theoretical foundations: Bayesian inference and the Kalman filter. Both derivations lead to the same uncertainty-weighted fusion expression:
$\mu_{f}=\frac{w_{i}\mu_{i}+w_{e}\mu_{e}}{w_{i}+w_{e}},$
with the inverse variance weights defined as $w_{m}=1/(\sigma_{m}^{2}+\epsilon)$ for Gaussian noise and $w_{m}=1/(\tilde{\sigma}_{m}^{2}+\epsilon)$ for the Student’s $t$ model. This dual theoretical basis justifies our approach, as it effectively leverages the inverse variances (or precisions) of the modality-specific estimates to account for their respective uncertainties, resulting in a robust and reliable fused representation for downstream tasks.
Appendix D Iterative Refinement of Fused State
After performing the sequential update to fuse the modalities over time using effective variances (i.e.,
$\tilde{\sigma}^{2}=\exp\Big{(}\log\sigma^{2}+\log\Big{(}\frac{\nu}{\nu+1}\Big{)}\Big{)}$
) to account for the heavy-tailed Student-T noise, small residual errors or microscopic uncertainties may still persist in the fused state. To address this, we introduce an iterative refinement step that further denoises and adjusts the fused representation.
The intuition behind this approach is similar to iterative error correction or gradient descent-based optimization: rather than relying solely on the initial fusion, we continuously refine the estimate to better capture the true latent state. Specifically, starting from the initial fused state $x_{\text{fusion}}^{0}$ obtained after the sequential update, we iteratively predict and subtract a residual correction. In each refinement step $i$ (for $i=0,1,\ldots,N-1$), a dedicated network $F(\cdot)$ takes the current state $x^{i}$ along with additional contextual information $c_{i}$ (which may include time-step context, current effective uncertainty estimates, and modality weights) and predicts a residual $\Delta x^{i}$:
$\Delta x^{i}=F(x^{i},c_{i}).$
This residual represents the remaining error in the current fused estimate. The state is then updated by subtracting a fraction of this residual, controlled by an attenuation parameter $\lambda_{i}$:
$x^{i+1}=x^{i}-\lambda_{i}\Delta x^{i}.$
After $N$ refinement steps, the final fused representation $x^{N}$ is obtained, which is expected to be more robust and accurate.
Theoretical Justification:
Even after uncertainty-weighted fusion and sequential updates—where effective variances derived via the Student-T model are used—the fused state may still contain imperfections due to noise, model approximation errors, or unmodeled dynamics. The refinement network is motivated by the following principles:
  • Residual Learning: The initial fused state $x_{\text{fusion}}^{0}$ is an approximation of the true latent state $y$, such that
    $y=x_{\text{fusion}}^{0}+\delta,$
    where $\delta$ is the residual error. The refinement network is designed to learn this residual:
    $F^{*}(x^{i},c_{i})\approx\mathbb{E}[y-x^{i}\mid x^{i},c_{i}],$
    so that the final estimate becomes
    $x^{N}=x_{\text{fusion}}^{0}-\sum_{i=0}^{N-1}\lambda_{i}F(x^{i},c_{i}).$
    This formulation is analogous to residual learning in deep networks, where modeling the error is often easier than directly predicting the target.
  • Diffusion Model Inspiration: Diffusion models iteratively denoise data by progressively removing noise from a corrupted input. Similarly, our iterative refinement can be viewed as a denoising process where each refinement step removes part of the residual error, thereby driving the fused state closer to the true latent representation.
  • Optimization Perspective: The refinement step can be interpreted as performing an additional optimization in function space. The subtraction of $\lambda_{i}\Delta x^{i}$ is akin to a gradient descent update that reduces an implicit error loss. Over multiple iterations, this results in a more accurate estimate, provided that the refinement network is properly designed and trained.
Empirical Benefits and Future Directions:
In practice, introducing a refinement network after the initial fusion offers several benefits:
  • Error Reduction: By learning the residual $\delta$, the final output $x^{N}$ achieves lower prediction error than the initial fused state.
  • Robustness: The iterative refinement is effective in mitigating the effects of heavy-tailed noise and unmodeled dynamics, leading to a more stable fused representation.
  • Enhanced Detail Recovery: Fine-grained details that might be lost in the initial fusion can be recovered through successive refinement, improving both quantitative metrics and qualitative performance.
Future work may explore:
  • Iterative or Recurrent Refinement: Extending the refinement process with additional iterative steps or a recurrent architecture that shares weights across iterations.
  • Uncertainty-Guided Refinement: Incorporating explicit uncertainty measures to guide the refinement network to focus on regions with high residual error.
  • Enhanced Loss Functions: Employing perceptual or adversarial losses in the refinement stage to better capture fine details and enhance the realism of the final output.
In summary, the iterative refinement network not only removes residual errors remaining after the initial uncertainty-weighted fusion but also draws strong inspiration from diffusion models’ denoising principles. This two-stage approach—first, a coarse fusion and then a fine, iterative refinement—provides a theoretically grounded and empirically validated method to enhance the final fused representation for downstream tasks.
Appendix E Derivation of KL Divergence
We present a detailed derivation of the KL divergence between two Gaussian distributions. Specifically, the KL divergence between a Gaussian distribution $\mathcal{N}(\mu_{m},\tilde{\sigma}_{m}^{2})$ and a standard normal distribution $\mathcal{N}(0,I)$ is derived in closed form as follows:
$KL\big{(}\mathcal{N}(\mu_{m},\tilde{\sigma}_{m}^{2})\|\mathcal{N}(0,I)\big{)}=\frac{1}{2}(\tilde{\sigma}_{m}^{2}+\mu_{m}^{2}-1-\log\tilde{\sigma}_{m}^{2}).$
The KL divergence between two probability distributions $p(x)$ and $q(x)$ is defined as:
$KL(p\|q)=\int p(x)\log\frac{p(x)}{q(x)}\,dx.$
Consider two Gaussian distributions:
$p(x)=\mathcal{N}(x;\mu_{m},\tilde{\sigma}_{m}^{2}),\quad q(x)=\mathcal{N}(x;0,1).$
Expanding explicitly, we obtain:
[…]
Evaluating the expectations under $p(x)$:
$\mathbb{E}_{p}[x]=\mu_{m},\quad\mathbb{E}_{p}[x^{2}]=\tilde{\sigma}_{m}^{2}+\mu_{m}^{2},\quad\mathbb{E}_{p}[(x-\mu_{m})^{2}]=\tilde{\sigma}_{m}^{2}.$
Substituting these expectations back into the expression gives:
$KL(p\|q)=\frac{1}{2}\log\frac{1}{\tilde{\sigma}_{m}^{2}}-\frac{1}{2\tilde{\sigma}_{m}^{2}}\tilde{\sigma}_{m}^{2}+\frac{1}{2}(\tilde{\sigma}_{m}^{2}+\mu_{m}^{2}).$
Simplifying further, we obtain the final closed-form expression:
$KL\big{(}\mathcal{N}(\mu_{m},\tilde{\sigma}_{m}^{2})\|\mathcal{N}(0,I)\big{)}=\frac{1}{2}(\tilde{\sigma}_{m}^{2}+\mu_{m}^{2}-1-\log\tilde{\sigma}_{m}^{2}).$
Appendix F Experiment Details
Experiments compute resources.
All experiments were conducted on a local workstation equipped with an AMD Ryzen Threadripper PRO 5955WX 16-Core Processor (32 threads) and a single NVIDIA RTX 6000 Ada Generation GPU (48GB VRAM). The system had 256GB of system RAM and 5GB VRAM and ran on Ubuntu 22.04. To improve training efficiency, we first precompute the video frame embeddings using the image and event encoders before training. This preprocessing step takes approximately 3 to 5 days. Once embeddings are extracted, training on the XD-Violence dataset takes around 3 hours, while training on the ShanghaiTech, UCF-Crime, and MSAD datasets completes within 1 hour.
F.1 Dataset Details
We evaluate our weakly supervised learning approach on four commonly used benchmark datasets for video anomaly detection: UCF-Crime [59], XD-Violence [66], ShanghaiTech [40], and MSAD [86]. UCF-Crime consists of 1,900 real surveillance videos, totaling 128 hours and covering 13 anomaly classes. XD-Violence comprises 4,754 untrimmed videos (217 hours), featuring 6 distinct anomalous or violent actions. ShanghaiTech includes 330 training and 107 test videos (approximately 317,000 frames), recorded in 13 scenes and labeled with 11 anomaly classes. MSAD features 720 videos from 14 different scenarios, annotated with 11 anomalies. Due to the data imbalance and the rarity of violent incidents in XD-Violence, we report the Average Precision (AP, %) to assess precision–recall balance, while for the other datasets, we measure performance using the Area Under the ROC Curve (AUROC, %).
Preprocessing
We employ a spatial augmentation strategy inspired by multi-crop evaluation. Specifically, each input video is first resized to a resolution of $224\times 224$, followed by the generation of 10 spatial crops: five fixed regions (top-left, top-right, bottom-left, bottom-right, and center) and their horizontally flipped counterparts. From each video, image frames are extracted and processed by the CLIP [54] (ViT-L/14) image encoder to obtain latent representations. For every 16 frames, we compute the average latent vector, denoted as $z_{x}$. In parallel, we generate synthetic events by computing pixel-wise changes between consecutive frames within each 16-frame segment using a threshold of 10/255 and a clamp value of 10. The resulting binary event maps are stacked and fed into the event encoder [31] (aligned with image encoder) to produce event representations $z_{e}$. We refer to these as "synthetic" since they are derived from RGB frames rather than captured by a true event sensor.
F.2 Implementation Details
Architecture Detail
Figure 1 illustrates the complete architecture used in our framework. Both the image and event encoders are implemented using the ViT-L/14 architecture from CLIP [54], with an embedding dimension of 768. The function $f$, responsible for modality-specific feature encoding, is implemented as a multi-layer attention module with 8 attention heads and 2 transformer layers. The functions $g$ and $h$, used to predict the mean and variance parameters for fusion, are each implemented as a single linear layer. The refinement network consists of a simple feedforward structure with a Linear–ReLU–Linear sequence to iteratively refine the fused representation.
Hyperparameters Detail
To reproduce the results reported in Table 1, we configure the model with the following hyperparameters: the degrees of freedom is set to $\nu{=}8$, the number of iterative refinement steps is $N{=}10$, and the uncertainty refinement weight is $\lambda_{r}{=}0.5$. The optimization uses the AdamW [43] optimizer with a learning rate of $2\times 10^{-5}$ and a batch size of 64 for 10 training epochs. We apply a MultiStepLR scheduler with milestones at epochs 4 and 8, and a decay factor of 0.1. The numerical stability term is set to $\epsilon{=}10^{-8}$. Additionally, for the regularization loss $\mathcal{L}_{\text{reg}}$, we fix both $\lambda_{1}$ and $\lambda_{2}$ to 0.5 throughout all experiments.
F.3 Results Detail
UCF-Crime [59] XD-Violence [66] ShanghaiTech [40] MSAD [86]
Class Image Event Fusion Class Image Event Fusion Class Image Event Fusion Class Image Event Fusion
Abuse 68.02 70.09 70.74 Fighting 79.59 67.81 84.76 Car 70.07 74.76 74.83 Assault 54.78 66.03 58.73
Arrest 72.21 47.09 75.05 Shooting 54.59 42.94 61.99 Chasing 94.49 84.36 91.33 Explosion 50.86 66.25 57.20
Arson 65.49 66.75 72.68 Riot 97.62 86.07 97.67 Fall 72.96 65.30 82.17 Fighting 71.74 79.75 81.14
Assault 56.44 72.03 72.58 Abuse 59.42 54.49 64.96 Fighting 76.91 63.48 83.64 Fire 71.97 49.44 71.23
Burglary 68.02 65.88 74.99 Car Accident 50.83 32.53 51.70 Monocycle 67.23 75.32 75.46 Object_falling 90.52 75.92 90.92
Explosion 56.33 57.64 63.46 Explosion 64.32 39.22 68.54 Robbery 76.74 87.26 90.43 People_falling 60.64 42.50 56.93
Fighting 58.14 79.27 62.81 Running 37.95 60.95 60.78 Robbery 68.10 66.90 70.95
RoadAccident 57.41 59.11 66.32 Skateboard 76.04 78.29 82.38 Shooting 71.20 86.87 77.88
Robbery 76.03 62.39 76.29 Throwing_object 89.63 83.14 91.95 Traffic_accident 62.23 70.08 70.93
Stealing 74.91 61.51 75.43 Vehicle 79.39 67.38 79.87 Vandalism 83.40 75.82 87.05
Shooting 60.95 38.42 62.26 Vaudeville 44.04 53.66 53.61 Water_incident 97.95 88.93 98.75
Shoplifting 64.27 73.29 85.72
Vandalism 66.89 63.05 69.23
AUC 86.77 78.67 89.13 AP 84.22 55.96 86.54 AUC 97.58 93.69 98.24 AUC 91.52 82.10 92.16
Ano AUC 66.56 63.94 72.49
Table 2: Per-class performance on UCF-Crime [59], XD-Violence [66], ShanghaiTech [40], and MSAD [86] for Figure 2.
Figure2 summarizes the per-class performance across four benchmark datasets: UCF-Crime, XD-Violence, ShanghaiTech, and MSAD. For each anomaly class, we report detection performance using the image modality, event modality, and their fusion. In most cases, the fusion consistently outperforms both unimodal inputs, highlighting the complementarity between image and event information. Notably, on UCF-Crime, classes such as Shoplifting and Arson show substantial gains from fusion. For XD-Violence, categories like Riot and Fighting exhibit strong improvements with fusion, despite relatively weak event-only performance. ShanghaiTech also shows a consistent pattern of fusion superiority across diverse scene types. In the MSAD dataset, fusion leads to higher detection scores for complex dynamic events such as Water_incident and Vandalism. These results emphasize the effectiveness of our uncertainty-guided multimodal fusion strategy in adapting to diverse scene contexts and anomaly types.
Appendix G Ablation Study
G.1 Sensitivity to Hyperparameter Settings
We conduct a comprehensive ablation study to examine the effect of key hyperparameters in our uncertainty-guided fusion framework. Specifically, we analyze the impact of the degrees of freedom $\nu$ in the Student-T distribution, the number of refinement steps $N$, the Laplace approximation precision $\epsilon$, and the refinement weight $\lambda_{r}$, all metrics is reported 10-times average with 1-standard deviation.
$\nu$ 2 4 6 8 10
AUC 87.98$\pm$0.33 88.34$\pm$0.30 88.38$\pm$0.40 88.67$\pm$0.45 88.11$\pm$0.47
Ano-AUC 70.06$\pm$0.72 70.69$\pm$0.72 70.88$\pm$0.86 71.50$\pm$1.02 70.22$\pm$0.86
N 0 10 20 30 40
AUC 86.77$\pm$0.18 88.67$\pm$0.45 88.41$\pm$0.43 88.48$\pm$0.28 88.60$\pm$0.31
Ano-AUC 67.41$\pm$0.43 71.50$\pm$1.02 71.27$\pm$0.61 71.36$\pm$0.59 71.45$\pm$0.58
$\epsilon$ $10^{-4}$ $10^{-6}$ $10^{-8}$
AUC 88.46$\pm$0.42 88.50$\pm$0.38 88.67$\pm$0.45
Ano-AUC 71.06$\pm$1.14 71.45$\pm$0.81 71.50$\pm$1.02
$\lambda_{r}$ 0.1 0.3 0.5 0.7 0.9
AUC 87.92$\pm$0.38 88.13$\pm$0.24 88.67$\pm$0.45 88.21$\pm$0.34 87.91$\pm$0.34
Ano-AUC 69.61$\pm$0.98 70.13$\pm$0.58 71.50$\pm$1.02 70.32$\pm$0.90 69.80$\pm$0.86
Table 3: Ablation study results on UCF-Crime with varying hyperparameters.
UCF-Crime
: Varying the degrees of freedom $\nu$ shows a gradual increase in both AUC and Ano-AUC scores, peaking at $\nu=8$. For the number of refinement steps $N$, performance increases significantly when $N$ changes from 0 to 10, and remains relatively stable for $N\geq 10$. The Laplace approximation precision $\epsilon$ results in only marginal differences across all values tested. In the case of the refinement weight $\lambda_{r}$, the best results are obtained at $\lambda_{r}=0.5$, while performance slightly degrades when the weight is set to more extreme values (0.1 or 0.9) Table 3.
Dataset Hyperparameter 2 / 0 / 1e-4 / 0.1 4 / 10 / – / 0.3 6 / 20 / 1e-6 / 0.5 8 / 30 / – / 0.7 10 / 40 / 1e-8 / 0.9
XD-Violence $\nu$ (AP) 87.31$\pm$0.77 87.39$\pm$0.65 87.03$\pm$0.52 87.63$\pm$0.54 86.64$\pm$0.51
N (AP) 86.01$\pm$1.10 87.63$\pm$0.54 87.52$\pm$0.40 87.71$\pm$0.70 87.54$\pm$0.60
$\epsilon$ (AP) 87.27$\pm$0.76 87.44$\pm$0.79 87.63$\pm$0.54
$\lambda_{r}$ (AP) 86.89$\pm$0.80 86.74$\pm$0.48 87.63$\pm$0.54 87.19$\pm$0.67 87.23$\pm$1.13
ShanghaiTech $\nu$ (AUC) 97.91$\pm$0.10 97.91$\pm$0.10 97.90$\pm$0.10 97.98$\pm$0.07 97.92$\pm$0.06
N (AUC) 97.94$\pm$0.08 97.98$\pm$0.07 97.90$\pm$0.12 97.90$\pm$0.11 97.93$\pm$0.07
$\epsilon$ (AUC) 97.91$\pm$0.12 97.90$\pm$0.08 97.98$\pm$0.07
$\lambda_{r}$ (AUC) 97.86$\pm$0.07 97.90$\pm$0.11 97.98$\pm$0.07 97.90$\pm$0.11 97.92$\pm$0.13
MSAD $\nu$ (AUC) 91.78$\pm$0.34 91.79$\pm$0.29 91.84$\pm$0.42 92.90$\pm$0.27 91.64$\pm$0.27
N (AUC) 90.69$\pm$0.35 92.90$\pm$0.27 92.29$\pm$0.30 92.72$\pm$0.29 92.69$\pm$0.21
$\epsilon$ (AUC) 91.77$\pm$0.27 91.81$\pm$0.33 92.90$\pm$0.27
$\lambda_{r}$ (AUC) 90.73$\pm$0.74 91.41$\pm$0.53 92.90$\pm$0.27 92.12$\pm$0.29 92.28$\pm$0.44
Table 4: Ablation study results across datasets (XD-Violence, ShanghaiTech, MSAD) under varying hyperparameters.
XD-Violence
: In the XD-Violence dataset, the degrees of freedom $\nu$ show stable performance across different settings, with AP values consistently around 87, peaking at 87.63 when $\nu=8$. Increasing the number of refinement steps $N$ leads to noticeable improvements, with AP rising from 86.01 at $N=2$ to 87.71 at $N=30$. The Laplace precision parameter $\epsilon$ exhibits minimal influence on performance, with AP differences within 0.36 across tested values. Adjusting the refinement weight $\lambda_{r}$ shows moderate effects, where the best AP of 87.63 is achieved at $\lambda_{r}=0.5$ Table 4.
ShanghaiTech
: For the ShanghaiTech dataset, the performance remains stable across different settings of the degrees of freedom $\nu$, with AUC scores ranging narrowly between 97.90 and 97.98. Varying the number of refinement steps $N$ does not lead to significant changes, though a slight increase is observed at $N=10$. The Laplace precision parameter $\epsilon$ shows minimal effect on AUC, with differences remaining within 0.08. Adjusting the refinement weight $\lambda_{r}$ yields the highest AUC of 97.98 at $\lambda_{r}=0.5$, while other values produce slightly lower but comparable performance Table 4.
MSAD
: In the MSAD dataset, increasing the degrees of freedom $\nu$ generally results in marginal improvements, peaking at $\nu=8$ with an AUC of 92.90. The number of refinement steps $N$ has a more pronounced impact, with AUC improving steadily from 90.69 at $N=0$ to 92.72 at $N=30$, then maintaining a similar level at $N=40$. The Laplace precision parameter $\epsilon$ again leads to only small variations, with values ranging from 91.77 to 92.90. For $\lambda_{r}$, AUC increases consistently as the parameter increases, with the highest value (92.28) observed at $\lambda_{r}=0.9$ Table 4.
G.2 Loss Configuration
Dataset Metric $\mathcal{L}_{\textbf{cls}}$ $\mathcal{L}_{\textbf{cls}}+\mathcal{L}_{\textbf{reg}}$ $\mathcal{L}_{\textbf{cls}}+\mathcal{L}_{\textbf{KL}}$ $\mathcal{L}_{\textbf{cls}}+\mathcal{L}_{\textbf{reg}}+\mathcal{L}_{\textbf{KL}}$
UCF-Crime AUC 88.04$\pm$0.28 87.78$\pm$0.46 88.24$\pm$0.19 88.67$\pm$0.45
Ano-AUC 69.79$\pm$0.85 69.73$\pm$1.00 70.46$\pm$0.69 71.50$\pm$1.02
XD-Violence AP 87.40$\pm$0.72 87.45$\pm$0.49 87.46$\pm$0.64 87.63$\pm$0.54
ShanghaiTech AUC 97.97$\pm$0.07 97.91$\pm$0.09 97.85$\pm$0.14 97.98$\pm$0.07
MSAD AUC 92.39$\pm$0.22 91.65$\pm$0.32 92.38$\pm$0.22 92.90$\pm$0.27
Table 5: Performance comparison under different loss configurations across datasets.
We assess the contribution of each loss component by evaluating four configurations: classification loss only ($\mathcal{L}_{\textbf{cls}}$), classification plus regularization ($\mathcal{L}_{\textbf{cls}}+\mathcal{L}_{\textbf{reg}}$), classification plus KL divergence ($\mathcal{L}_{\textbf{cls}}+\mathcal{L}_{\textbf{KL}}$), and the full combination ($\mathcal{L}_{\textbf{cls}}+\mathcal{L}_{\textbf{reg}}+\mathcal{L}_{\textbf{KL}}$) across UCF-Crime, XD-Violence, ShanghaiTech, and MSAD datasets (Table 5).
In UCF-Crime, adding the KL loss improves AUC from 88.04% to 88.24% and Ano-AUC from 69.79% to 70.46%, while the full combination further boosts AUC and Ano-AUC to 88.67% and 71.50%, respectively. In XD-Violence, ShanghaiTech, and MSAD, performance remains largely stable across settings, with slight improvements under the full loss configuration. In particular, MSAD AUC increases from 92.39% to 92.90% with the full loss. Overall, adding $\mathcal{L}_{\textbf{KL}}$ consistently yields benefits, while the effect of $\mathcal{L}_{\textbf{reg}}$ alone is minor. The full configuration achieves the best or comparable results across all benchmarks.
Appendix H Fusion Details
To evaluate whether the proposed framework performs as intended, we analyze the behavior of the uncertainty weights $w_{m}$ (Eq. 4) when one modality is partially corrupted. Specifically, we aim to test whether the model’s dynamic uncertainty estimation mechanism can correctly respond to degraded sensor input. Rather than injecting additive noise—which may lead to complex and unpredictable interactions within attention-based encoders—we opt for a controlled masking strategy. In attention networks, even small perturbations can propagate nonlinearly across dimensions, making it difficult to interpret the resulting change in uncertainty due to entangled feature dependencies. Additionally, adversarial perturbations rely on gradients computed after the refinement stage, making it difficult to isolate the direct effect on modality-specific uncertainty. In contrast, masking fixed proportions of input features allows us to deterministically degrade the modality in a localized and interpretable manner, providing a clean testbed for evaluating the reliability and sensitivity of uncertainty estimation.
To assess the reliability of our uncertainty-weighted fusion mechanism, we conduct controlled modality-specific perturbation experiments across four datasets. We simulate degradation by randomly masking a proportion $\rho\in\{0.05,0.10,0.20,0.30,0.50\}$ of the latent feature dimensions in either the image modality ($z_{x}$) or the event modality ($z_{e}$). All experiments report standard performance metrics, including AUC or AP for detection quality, Brier score for probabilistic calibration, and KL divergence to quantify the shift between predictions made on clean inputs and those made under masking. Additionally, we track uncertainty weights $w_{x}$ and $w_{e}$ for both modalities, including breakdowns on abnormal and normal video segments, to understand how the model reallocates modality-level confidence under degradation.
Noise Type Masked Level AUC (%) Brier KL $\Delta w_{e}$ $\Delta w_{e}^{\mathrm{ab}}$ $\Delta w_{e}^{\mathrm{n}}$ $\Delta w_{x}$ $\Delta w_{x}^{\mathrm{ab}}$ $\Delta w_{x}^{\mathrm{n}}$
CLEAN 0 89.09 0.1205 0.0000 0.4760 0.4744 0.4761 0.5240 0.5256 0.5239
EV_NOISE 0.05 88.92 0.1238 0.0032 0.4757 0.4742 0.4758 0.5243 0.5258 0.5242
0.10 88.80 0.1249 0.0062 0.4756 0.4742 0.4758 0.5244 0.5258 0.5242
0.20 88.68 0.1252 0.0104 0.4757 0.4743 0.4758 0.5243 0.5257 0.5242
0.30 88.64 0.1238 0.0133 0.4757 0.4743 0.4758 0.5243 0.5257 0.5242
0.50 88.35 0.1216 0.0198 0.4758 0.4745 0.4759 0.5242 0.5255 0.5241
IMG_NOISE 0.05 88.57 0.1261 0.0258 0.4764 0.4749 0.4766 0.5236 0.5251 0.5234
0.10 87.90 0.1295 0.0548 0.4769 0.4755 0.4771 0.5231 0.5245 0.5229
0.20 87.05 0.1290 0.1066 0.4779 0.4766 0.4780 0.5221 0.5234 0.5220
0.30 86.57 0.1238 0.1445 0.4788 0.4776 0.4790 0.5212 0.5224 0.5210
0.50 85.74 0.1097 0.2021 0.4808 0.4798 0.4809 0.5192 0.5202 0.5191
Table 6: Fusion metrics on the UCF-Crime dataset under varying noise settings. AUC is reported as a percentage.
Noise Type Masked Level AP (%) Brier KL $\Delta w_{e}$ $\Delta w_{e}^{\mathrm{ab}}$ $\Delta w_{e}^{\mathrm{n}}$ $\Delta w_{x}$ $\Delta w_{x}^{\mathrm{ab}}$ $\Delta w_{x}^{\mathrm{n}}$
CLEAN 0 88.26 0.0735 0.0000 0.4661 0.4623 0.4672 0.5339 0.5377 0.5328
EV_NOISE 0.05 88.12 0.0741 0.0020 0.4661 0.4623 0.4672 0.5339 0.5377 0.5328
0.10 88.00 0.0747 0.0040 0.4661 0.4623 0.4671 0.5339 0.5377 0.5329
0.20 87.74 0.0761 0.0083 0.4660 0.4624 0.4671 0.5340 0.5376 0.5329
0.30 87.51 0.0769 0.0121 0.4660 0.4624 0.4670 0.5340 0.5376 0.5330
0.50 87.13 0.0792 0.0194 0.4658 0.4623 0.4668 0.5342 0.5377 0.5332
IMG_NOISE 0.05 87.95 0.0794 0.0267 0.4665 0.4628 0.4676 0.5335 0.5372 0.5324
0.10 87.64 0.0852 0.0546 0.4670 0.4633 0.4680 0.5330 0.5367 0.5320
0.20 86.60 0.0986 0.1162 0.4679 0.4644 0.4689 0.5321 0.5356 0.5311
0.30 85.86 0.1109 0.1788 0.4687 0.4655 0.4697 0.5313 0.5345 0.5303
0.50 84.23 0.1366 0.3082 0.4705 0.4677 0.4714 0.5295 0.5323 0.5286
Table 7: Fusion metrics on the XD-Violence dataset under varying noise settings.
Noise Type Masked Level AUC (%) Brier KL $\Delta w_{e}$ $\Delta w_{e}^{\mathrm{ab}}$ $\Delta w_{e}^{\mathrm{n}}$ $\Delta w_{x}$ $\Delta w_{x}^{\mathrm{ab}}$ $\Delta w_{x}^{\mathrm{n}}$
CLEAN 0 98.17 0.0402 0.0000 0.4718 0.4633 0.4723 0.5282 0.5367 0.5277
EV_NOISE 0.05 98.13 0.0401 0.0006 0.4720 0.4634 0.4725 0.5280 0.5366 0.5275
0.10 98.11 0.0398 0.0022 0.4722 0.4635 0.4727 0.5278 0.5365 0.5273
0.20 98.07 0.0399 0.0038 0.4726 0.4638 0.4731 0.5274 0.5362 0.5269
0.30 98.06 0.0393 0.0053 0.4729 0.4639 0.4734 0.5271 0.5361 0.5266
0.50 98.06 0.0383 0.0082 0.4734 0.4641 0.4739 0.5266 0.5359 0.5261
IMG_NOISE 0.05 97.66 0.0430 0.0162 0.4726 0.4647 0.4731 0.5274 0.5353 0.5269
0.10 97.54 0.0434 0.0308 0.4733 0.4651 0.4738 0.5267 0.5349 0.5262
0.20 96.29 0.0477 0.0758 0.4750 0.4678 0.4754 0.5250 0.5322 0.5246
0.30 95.09 0.0520 0.1084 0.4764 0.4696 0.4768 0.5236 0.5304 0.5232
0.50 90.72 0.0611 0.2180 0.4801 0.4754 0.4803 0.5199 0.5246 0.5197
Table 8: Fusion metrics on the ShanghaiTech dataset under varying noise settings.
Noise Type Masked Level AUC (%) Brier KL $\Delta w_{e}$ $\Delta w_{e}^{\mathrm{ab}}$ $\Delta w_{e}^{\mathrm{n}}$ $\Delta w_{x}$ $\Delta w_{x}^{\mathrm{ab}}$ $\Delta w_{x}^{\mathrm{n}}$
CLEAN 0 92.39 0.1119 0.0000 0.4807 0.4786 0.4814 0.5193 0.5214 0.5186
EV_NOISE 0.05 92.27 0.1126 0.0024 0.4808 0.4787 0.4815 0.5192 0.5213 0.5185
0.10 92.19 0.1129 0.0048 0.4810 0.4788 0.4816 0.5190 0.5212 0.5184
0.20 91.94 0.1149 0.0088 0.4812 0.4791 0.4819 0.5188 0.5209 0.5181
0.30 91.89 0.1153 0.0116 0.4814 0.4791 0.4821 0.5186 0.5209 0.5179
0.50 91.86 0.1168 0.0158 0.4817 0.4794 0.4824 0.5183 0.5206 0.5176
IMG_NOISE 0.05 92.27 0.1117 0.0066 0.4811 0.4790 0.4817 0.5189 0.5210 0.5183
0.10 92.10 0.1118 0.0125 0.4816 0.4795 0.4822 0.5184 0.5205 0.5178
0.20 91.86 0.1111 0.0220 0.4824 0.4804 0.4830 0.5176 0.5196 0.5170
0.30 91.63 0.1116 0.0303 0.4832 0.4813 0.4838 0.5168 0.5187 0.5162
0.50 91.52 0.1106 0.0425 0.4848 0.4827 0.4854 0.5152 0.5173 0.5146
Table 9: Fusion metrics on the MSAD dataset under varying noise settings.
Across all four datasets, our method consistently demonstrates robust uncertainty-guided fusion behavior. When degradation is applied to the event modality ($z_{e}$), performance remains stable—AUC or AP typically drops by less than 1%, and uncertainty weights show minimal change. This suggests that the model is not overly sensitive to event corruption and maintains reliable fusion under partial degradation.
In contrast, masking the image modality ($z_{x}$) produces more pronounced effects, especially on appearance-dependent datasets such as UCF-Crime and XD-Violence. At the highest masking level, UCF-Crime experiences a 2.35% drop in AUC, and XD-Violence shows nearly a 5% drop in AP, accompanied by KL divergence increases up to 0.3082. Uncertainty weights reflect this asymmetry: $w_{x}$ consistently decreases while $w_{e}$ increases, indicating that the model dynamically downweights unreliable image features and reallocates confidence toward the event modality.
On datasets with higher modality redundancy—such as ShanghaiTech and MSAD—both performance and uncertainty remain relatively stable under perturbation. ShanghaiTech maintains an AUC above 98% under event noise and above 90% under severe image masking, while MSAD exhibits only minor fluctuations across all metrics. This confirms that the model can maintain balanced modality fusion when both modalities offer sufficient information.
While the average change in uncertainty weights ($\Delta w_{m}$) is numerically small—typically below 1%—this is largely due to aggregation over all vector indices and time steps. A finer-grained analysis reveals that individual latent dimensions can shift by as much as 30% under value-level masking, demonstrating substantial feature-wise modulation. Moreover, uncertainty reallocation is more pronounced in abnormal segments compared to normal ones, indicating that the model adapts more sensitively in semantically critical regions. Finally, increases in Brier score under corruption reflect growing misalignment between predicted probabilities and ground truth labels, particularly under image degradation, further confirming that the model’s confidence dynamically adjusts in response to input quality.
Appendix I Visualizatation of Anomaly Detection with Uncertainty

### v2

@@
 Most existing video anomaly detectors rely solely on RGB frames, which lack the temporal resolution needed to capture abrupt or transient motion cues—key indicators of anomalous events.
-To address this limitation, we propose Image-Event Fusion for Video Anomaly Detection (IEF-VAD), a framework that synthesizes event representations directly from RGB videos and fuses them with image features through a principled, uncertainty-aware process.
-The system (i) models heavy-tailed sensor noise with a Student’s-t likelihood, deriving value-level inverse-variance weights via a Laplace approximation; (ii) applies Kalman-style frame-wise updates to balance modalities over time; and (iii) iteratively refines the fused latent state to erase residual cross-modal noise.
+To address this limitation, we propose Image-Event Fusion for Video Anomaly Detection (IEF-VAD), a framework that synthesizes event directly from RGB videos and fuses them through a principled, uncertainty-aware process.
+The system (i) models heavy-tailed sensor noise with a Student’s t likelihood, deriving value-level inverse-variance weights via a Laplace approximation; (ii) applies Kalman-style frame-wise updates to recalibrate modality weights based on temporal context; and (iii) iteratively refines the fused latent state to remove residual cross-modal noise.
 Without any dedicated event sensor or frame-level labels, IEF-VAD sets a new state of the art across multiple real-world anomaly detection benchmarks.
-These findings highlight the utility of synthetic event representations in emphasizing motion cues that are often underrepresented in RGB frames, enabling accurate and robust video understanding across diverse applications without requiring dedicated event sensors.
-Code and models are available at https://github.com/EavnJeong/IEF-VAD.
+These findings highlight the utility of synthetic event in emphasizing motion cues that are often underrepresented in RGB frames, enabling accurate and robust video understanding across diverse applications without requiring dedicated event sensors.
+The code is available at: https://github.com/EavnJeong/IEF-VAD
 1 Introduction
@@
 5 Conclusion
-IEF-VAD proves that motion-centric synthetic event streams distilled from ordinary videos can be fused with RGB features to push video-anomaly detection beyond the limits of frame-based models.
-The approach establishes new state-of-the-art scores on UCF-Crime, XD-Violence, ShanghaiTech, and MSAD, while a value-level masking study shows its precision weights shift adaptively across latent dimensions—making the fusion logic transparent and reliable.
-We expect these findings to catalyse broader use of synthetic event data and to drive future work on uncertainty-aware multimodal learning for real-time surveillance.
+IEF-VAD proves that motion-centric synthetic event streams distilled from ordinary videos can be fused with RGB frame to push video-anomaly detection beyond the limits of frame-based models.
+The approach establishes new state-of-the-art scores on UCF-Crime, XD-Violence, ShanghaiTech, and MSAD, while a value-level masking study shows its weights shift adaptively across latent dimensions, revealing a reliable reliance on modality-specific cues.
+We expect these findings to catalyse broader use of synthetic event data and demonstrate its potential for principled, effective multimodal fusion in video understanding.
 Limitations and Future Work.
@@
 First, approximating each modality’s noise with a diagonal covariance neglects cross-feature correlations; future work could adopt lightweight structured or low-rank covariance estimators to enrich value-level fusion.
-Second, the fixed regularization weights $\lambda_{1,2}$ and Student-t degrees-of-freedom $\nu$ limit adaptability; jointly learning or meta-learning these parameters would enable the model to recalibrate trust as modality quality shifts.
+Second, the fixed regularization weights $\lambda_{1,2}$ and Student-t degrees-of-freedom $\nu$ limit adaptability; learning these parameters jointly with the model, or via meta-learning, would allow the system to adjust its reliance on each modality based on their varying quality.
 Third, replacing simple weight averaging with uncertainty-conditioned cross-modal attention, hierarchical gating, or iterative message passing may yield a more expressive fusion mechanism that leverages inter-modal disagreement as informative signal, paving the way for stronger anomaly detection and broader multimodal perception capabilities.